Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

confusing gmatclub probability

Expert replies
by arora007 » Tue Jul 06, 2010 9:40 pm
What is the probability that one of the two integers randomly selected from range 20-29 is prime and the other is a multiple of 3?
(The numbers are selected independently of each other, i.e. they can be equal)



a) 0.06
b) 0.12
c) 0.15
d) 0.18
e) 0.20

[spoiler]
OA is b , but how?? i do not understand a part of the explanation.

Prime integers: 23 and 29. Multiples of 3: 21, 24, and 27.
The probability that the first number is prime while the second is a multiple of 3 = (2/10)*(3/10) = 0.06. [should this not be (2/10)*(3/9) ]
The probability that the first number is a multiple of 3 while the second is prime = (3/10)*(2/10) = 0.06. [should this not be (3/10)*(2/9) ]
The probability that one of the two integers is prime and the other is a multiple of 3 = 0.06 + 0.06 = 0.12 .
[and the resulting sum of probabilities= 2/15 ??][/spoiler]
https://www.skiponemeal.org/
https://twitter.com/skiponemeal
Few things are impossible to diligence & skill.Great works are performed not by strength,but by perseverance

pm me if you find junk/spam/abusive language, Lets keep our community clean!!
Join the discussion
Source: — Problem Solving |

by kvcpk » Tue Jul 06, 2010 10:18 pm
Prime numbers are 23,29
Multiple of 3 are 21,24,27

Now, probability of picking prime is 2/10
probability of picking mult of 3 is 3/10

probability of both can be done together is 3/10 * 2/10 = 6/100

But the order can change. so 2*6/100 = 12/100 = 0.12

Hope this helps!!
Join the discussion

by arora007 » Tue Jul 06, 2010 10:29 pm
But when i select either of the prime no. or the multiple of three.. wouldn't 9 be left?

so

2/10 * 3/9 and not 2/10 * 3/10

is it that i am not getting the message for the text ??

"The numbers are selected independently of each other, i.e. they can be equal"
https://www.skiponemeal.org/
https://twitter.com/skiponemeal
Few things are impossible to diligence & skill.Great works are performed not by strength,but by perseverance

pm me if you find junk/spam/abusive language, Lets keep our community clean!!
Join the discussion

by kvcpk » Tue Jul 06, 2010 11:32 pm
arora007 wrote:But when i select either of the prime no. or the multiple of three.. wouldn't 9 be left?

so

2/10 * 3/9 and not 2/10 * 3/10

is it that i am not getting the message for the text ??

"The numbers are selected independently of each other, i.e. they can be equal"
yupp.. numbers are chosen independently.. so you need not change the count.
Join the discussion

by immaculatesahai » Fri Dec 16, 2011 2:19 am
@arora007,

I agree with your reasoning. The question "must" mention that " the numbers are replaceable" in order to avoid confusion. But your reasoning is correct.

The question is quite poor. Would suggest that you stay off poorly constructed questions such as these.

Enjoy !!!
Join the discussion